ABC Ltd is evaluating to invest in two projects. Project X may yield a return of £1.5m with a probability of 20%, or a return of £5m with a probability of 60%. Project Y, instead, may earn a negative return of £3m with a probability of 70% or a positive return of £7m with a probability of 30%.

1) How much is the expected return for project X and Y?


1
Expert's answer
2021-07-08T14:35:17-0400

Solution:

Expected return = "\sum"(Return i "\times" Probability i)


Expected return for project X = "\sum"(1,500,000 "\times" 0.2) + (5,000,000 "\times" 0.60) = 300,000 + 3000,000 = 3,300,000

Expected return for project X = £3,300,000


Expected return for project Y = "\sum"(-3000,000 "\times" 0.7) + (7,000,000 "\times" 0.30) = -2,100,000 + 2,100,000 = 0

Expected return for project X = £0